Given: 4 NO_2(g) + O_2(g) rightarrow 2 N_2O_5(g) d
Answer #1

9.

Given for,

4NO2 + O2 --> 2N2O5 dHb = -110.2 kJ

Divide this equation by 2, and invert the reaction,

fo reverse reaction,

N2O5 <==> 2NO2 + 1/2O2

dH = 110.2/2 = +55.1 kJ

11. moles of C = 49.97/12 = 4.16 mol

moles of H = 10.51/1 = 10.51 mol

moles of F = 39.52/19 = 2.08 mol

divide by smallest number

C = 4.16/2.08 = 2

H = 10.51/2.08 = 5

F = 2.08/2.08 = 1

Empirical formula = C2H5F